The upper boundaries for a distribution of waiting times (in seconds) in a grocery store aisle are 47, 60, 73, and 86. List the value for each lower boundary in this distribution
Answer #1

ANSWER:

Given that,

Upper boundaries are 47, 60, 73, 86

Class width = 60-47-1= 13-1=12

So lower boundaries are 47-12= 35

60-12= 48

73-12= 61

86-12= 74

So lower boundaries are 35, 48, 61, 74